OpenStudy (twistrose):

Solve for x. 9x=3x−36 −6−6 −4−4 4 6

OpenStudy (quickstudent):

You put '=' twice in the equation. Did you make a typo?

OpenStudy (twistrose):

yes im sorry hold on

OpenStudy (quickstudent):

OK, here's how you solve for x here. 9x = 3x − 36 9x - 3x = 3x - 36 - 3x 6x = -36 6x/6 = -36/6 x = -6

OpenStudy (otherworldly):

^^^^^ ur too fast for me XD

OpenStudy (twistrose):

XD

OpenStudy (twistrose):

Thanks you both ya'll's help

OpenStudy (quickstudent):

XD @OtherWorldly

OpenStudy (otherworldly):

XP